A steam engine (assume a Carnot engine ) has an efficiency of 34.8%.
If the waste heat has a temperature of 76◦C, what is the temperature of the boiler?
Answer in units of ◦C.
1
Expert's answer
2019-04-30T09:44:45-0400

The efficiency of a Carnot engine:


"\eta=1- \frac{T_C}{T_H}=1- \frac{t_C+273}{t_H+273}"

"0.348=1- \frac{76+273}{t_H+273}"

"t_H=262^\circ C"